### v2

# An Efficient Intelligent Semi-Automated Warehouse Inventory Stocktaking System

Abstract
As supply chain management continues to evolve, efficient inventory management systems have become increasingly crucial. However, traditional manual methods often struggle to meet the complexities of modern market demands, particularly when it comes to data accuracy, delays in monitoring, and the heavy reliance on subjective experience for forecasting. This study introduces an intelligent, semi-automated barcode-based inventory management system designed to overcome these challenges. The system integrates barcode technology with a distributed architecture, combined with big data analytics and machine learning for real-time tracking and accurate inventory predictions Its performance has been validated through multiple simulation tests, where it has outperformed traditional RFID technology in certain cases. Through careful system design, technology exploration, and validation, this research demonstrates the significant potential of this intelligent system in improving inventory management efficiency and accuracy.
Keywords stocktaking $\cdot$ inventory  $\cdot$ stock count  $\cdot$ RFID  $\cdot$ semi-automated  $\cdot$ warehouse
1 Introduction
In recent years, fully automated inventory technologies have experienced significant advancements, with the utilization of next-generation ultra-high-frequency RFID (Radio-Frequency Identification) employing radio waves for real-time identification and tracking of tagged objects. The application of RFID in warehouse management traces back to the 1990s, with initiatives by major players such as Walmart and the US Department of Defense. However, it wasn’t until the 2000s that RFID technology gained widespread adoption due to the establishment of standards and cost reductions [13]. While new RFID systems have demonstrated impressive accuracy in real-time inventory monitoring and complex tasks like expiration date tracking, the full extent of their potential remains unpredictable. This raises questions about the limited adoption of these powerful systems in large manufacturing factories.
Although fully automated inventory technology is promising, warehouse management is not yet ready for full adoption.. Prior research, such as Dong et al. [20], has highlighted a fundamental issue in large-scale RFID deployments: the complexity of load balancing in dense inventory environments. Their analysis demonstrated that minimizing RFID reader overload is an NP-hard problem, meaning that maintaining consistent performance in high-density warehouses is computationally intensive and inherently inefficient. This challenge is further exacerbated by environmental factors—Griffiths et al. [8] found that RFID accuracy significantly deteriorates when tags are placed near metals or liquids, making RFID unreliable in diverse warehouse conditions.
Beyond technical limitations, RFID adoption also faces economic and operational barriers. Huber et al. [7] identified that the high cost of RFID tags, readers, and integration with existing warehouse management systems creates a financial burden for businesses. While RFID promises real-time stock tracking, achieving full automation often requires substantial infrastructure investment, which many companies find prohibitive. Furthermore, studies such as White et al. [15] indicate that fully automated RFID inventory systems may not be as cost-effective as initially anticipated, due to ongoing maintenance expenses and frequent calibration requirements.
Due to these limitations, it is clear that a fully automated RFID-based inventory management system is not always the optimal solution, particularly for warehouses transitioning from traditional manual stocktaking methods. The gap between RFID’s potential and its real-world feasibility necessitates an alternative approach that balances automation with operational practicality. This study proposes a semi-automated inventory system that leverages barcode scanning and real-time data analytics to bridge this gap. Unlike RFID, which suffers from environmental interference and high implementation costs, barcode-based inventory tracking provides a more stable, cost-effective, and adaptable alternative.
The proposed semi-automated inventory system integrates handheld barcode scanning terminals with an intelligent backend architecture, ensuring real-time ERP synchronization and dynamic stock management. The system follows a client-server model, where Android-based PDA devices capture barcode data and communicate with the backend via a RESTful API. The backend, running on an Ubuntu Linux environment with a distributed Hadoop HDFS database, cross-checks scanned records against SAP S/4HANA ERP for real-time stock validation.
A key innovation of this system is integrating Apache Flink’s watermarking mechanism, which enables real-time tracking of inbound and outbound inventory. By leveraging event-time processing, Flink dynamically updates inventory levels, ensuring high accuracy even in high-throughput warehouse environments. This approach effectively mitigates latency issues in traditional stock reconciliation, allowing instant discrepancy detection.
To improve anomaly detection and stock integrity, the backend uses AI-driven analytics with the Isolation Forest algorithm. It converts Bin, Batch, and Handling Unit (HU) data into feature vectors and assigns anomaly scores to identify discrepancies.
This hierarchical anomaly detection method fully utilizes structured stocktaking data to improve inventory accuracy, enable real-time monitoring, and prevent errors proactively. Combined with scalable big data processing, it provides a cost-effective, flexible alternative to RFID-based solutions, making it highly suitable for modern warehouse management and digital transformation.
By critically assessing RFID’s limitations and offering a structured alternative, this research provides a clear pathway for businesses seeking to modernize their inventory management without the financial and operational burdens associated with full RFID adoption. This study contributes to the field by presenting a scalable and pragmatic inventory solution, offering valuable insights for warehouse managers and supply chain professionals navigating the transition toward digitalized stock control.
2 Related Work and Research Gap
Barcode-based inventory management systems have long been a staple in warehouse operations due to their affordability and simplicity. However, despite their widespread use, they exhibit several limitations, especially in large-scale or high-turnover environments. One major issue with traditional barcode systems is the reliance on batch synchronization for inventory updates. As a result, data is often outdated by the time it is processed, which hinders timely decision-making. The manual scanning process, though essential for the accuracy of the data, introduces delays as updates occur only once the scanning process is complete.
The dependence on human labor also opens the door to various types of errors. Barcode scanning systems are still heavily reliant on operators manually scanning each item, which leads to mistakes such as missed scans, misidentification of items, or duplication of scans. Such errors accumulate over time, significantly affecting inventory accuracy and requiring additional manual checks to resolve discrepancies. As warehouses grow, these inefficiencies become more pronounced, limiting the scalability and operational effectiveness of barcode-based systems.
Existing barcode systems also lack intelligent anomaly detection mechanisms. Discrepancies in inventory data are often not identified until manually flagged by operators, which can delay responses to issues like stockouts or overstocking. This lack of real-time anomaly detection exacerbates operational inefficiencies, particularly in large-scale warehouses where the time taken to identify and correct discrepancies is critical.
2.1 Semi-Automated Barcode-Based Systems
Semi-automated barcode-based systems, while still reliant on manual scanning, have been increasingly integrated with software systems that streamline the inventory process. Research has shown that integrating barcode scanning with cloud-based platforms can help improve data accuracy and enable real-time synchronization between the warehouse and ERP systems.An study addresses the inventory management issues of the Zambia Air Force (ZAF) by proposing an automated inventory management system based on cloud architecture and barcode technology to resolve problems caused by manual operations, such as inventory errors, inefficient management, item losses, and high costs [21]. System testing demonstrated that barcode scanning improved inventory accuracy by 95%, was over three times faster than manual management, and enhanced security through CCTV monitoring. However, the study has certain limitations, including the lack of consideration for more advanced technologies such as RFID, insufficient security analysis, unquantified CCTV effectiveness, and the absence of a return on investment (ROI) assessment.
2.2 RFID-Based Semi-Automated Inventory Counting
Similarly, RFID represents the latest technology in next-generation warehouse inventory management and has been included in comparative studies to evaluate its advantages and limitations in improving inventory management efficiency. A study [15] conducted an experimental comparison between barcoding and radio frequency identification (RFID) technologies in inventory management, focusing on scanning speed, accuracy, equipment failure rate, and applicable environments, with testing carried out in a cold chain warehouse. The results showed that RFID scanning speed was 2.5 times faster than barcoding, but its error rate (46.5%) and equipment failure rate (37%) were higher, and its performance in cold chain environments was unstable. The study also highlighted that two-dimensional barcodes (2D Barcode) are narrowing the performance gap with RFID and suggested that future research should explore hybrid systems (Hybrid Systems) that integrate the advantages of both technologies. However, the study lacked detailed cost analysis and long-term return on investment (ROI) evaluation and was limited to cold chain warehouses, without addressing the needs of other industries. Future advancements in expanding the scope of experiments, optimizing RFID equipment, researching hybrid systems, and analyzing its potential in supply chain big data management could enhance the intelligence and efficiency of inventory management. However, while these solutions address some of the shortcomings of traditional barcode systems, they still rely heavily on manual scanning and lack intelligent anomaly detection capabilities, leaving them vulnerable to human error and delayed issue resolution.
In session 2.21, Dong’s research revealed that load balancing problem in large-scale RFID systems reveals that RFID readers in high-density tag environments are prone to uneven load distribution, data loss, and excessive energy consumption. However, the study does not explore dynamic environments, cost analysis, and security issues in depth, which share certain similarities with the conclusions drawn by White et al. (2007) in their comparison of barcoding and RFID technologies [15], which also served as one of the motivations for this research.
RFID technology consists of three main components:
  • RFID Tags: Embedded microchips attached to inventory items for unique identification.
  • RFID Readers: Devices that emit radio waves to activate and read RFID tags.
  • RFID Middleware: A software layer responsible for processing tag data and integrating with warehouse management systems (WMS) or enterprise resource planning (ERP) platforms.
  1. 1.
    Readability Issues in Complex Warehouse Environments
    Signal interference from metal surfaces, liquids, and electromagnetic interference (EMI) significantly affects scanning accuracy. The probability of a successful RFID read $P_{\text{read}}$ decreases with increasing tag density:
    $P_{\text{read}}=\begin{cases}1-\frac{N_{\text{tags}}}{N_{\text{max}}},&\text{if }N_{\text{tags}}\leq N_{\text{max}},\\ 0,&\text{if }N_{\text{tags}}>N_{\text{max}}.\end{cases}$ (1)
    where $N_{\text{tags}}$ is the number of RFID tags in the scanning zone, and $N_{\text{max}}$ is the reader’s tag capacity before data loss occurs.
  2. 2.
    High Cost of Implementation
    RFID deployment involves substantial infrastructure investment, including RFID readers, middleware, and ERP integration. Tagging costs remain high, especially for environments requiring metal-compatible RFID tags.
  3. 3.
    Network Congestion and Data Latency
    Large-scale RFID deployments produce high-frequency data transmissions, causing network congestion and data loss:
    $P_{\text{loss}}=1-\frac{C_{\text{max}}}{M\times R}$ (2)
    where $C_{\text{max}}$ is server processing capacity, $M$ is the number of RFID readers, and $R$ is the data rate per reader.
  4. 4.
    Complex Integration with Existing Systems
    Unlike barcode-based solutions, RFID requires specialized ERP connectors for real-time data processing. Many legacy warehouse management systems lack native RFID support, necessitating additional middleware.
These challenges necessitate an alternative inventory tracking approach that balances automation, cost-effectiveness, and ease of deployment.
2.3 Research Gap
Current research in semi-automated inventory management often focuses on RFID as the primary solution. However, there is a growing need for alternative approaches that are cost-effective, scalable, and easy to deploy. Existing studies on barcode-based systems have explored their integration with cloud computing and mobile devices [21], but these systems still fall short in addressing the need for real-time updates and intelligent anomaly detection.
To fill this gap, our system combines the best of barcode technology with cutting-edge stream processing and AI capabilities, offering a more balanced solution that minimizes costs, reduces errors, and scales efficiently.
We compare our system to RFID-based, traditional manual inventory counting, and barcode-based systems in the following table:
Table 1: Comparison of Inventory Management Approaches
Feature RFID-Based System Traditional Manual Stocktaking Proposed System (Flutter + Flink + AI)
Cost High (\$500K+ for full deployment) Low (Minimal investment) Low (Standard barcode scanners + PDA)
Stocktaking Speed Fast (but affected by signal interference) Slow (Manual item-by-item scanning) Fast (Automated barcode scanning + AI anomaly detection)
Error Rate Moderate (RFID misreads, environmental factors) High (Human errors) Low (AI-assisted detection, automated validation)
Real-time Sync Requires additional ERP integration No real-time synchronization Instant updates via Flink stream processing
Scalability Requires expensive infrastructure Limited (Depends on workforce availability) High (Easily adaptable)
Environmental Reliability Prone to interference (Metal/liquids) Not affected Not affected (Barcode scanning is robust)
Ease of Use Requires training Time-consuming User-friendly with AI-assisted alerts
Operational Efficiency Medium (Some automation, but requires ERP adjustments) Low (Labor-intensive, slow updates) High (AI minimizes manual intervention, increases efficiency)
Maintenance Cost High (Frequent tag replacements, system tuning) Low (Minimal maintenance) Low (Barcode-based, minimal upkeep)
Implementation Complexity High (Complex hardware/software integration) Low (No complex infrastructure required) Low (Standard barcode scanning + lightweight cloud integration)
Training Time for Operators 6+ hours (Needs training in RFID usage and ERP) 3+ hours (Manual stocktaking procedures) 1-2 hours (Simple barcode scanning with AI-assisted alerts)
Overall Cost-Effectiveness Expensive (High initial cost, maintenance, and operational expenses) Low-cost (But inefficient in large-scale operations) Optimal balance of cost, speed, and accuracy
3 System Overview
Figure 1: System Architecture Diagram
Figure 1 illustrates the overall architecture of the developed system. The system architecture consists of four core components. First, handheld PDA terminals running a Flutter-based inventory application interact with the backend through a REST API. Second, a distributed processing backend utilizes Apache Flink for event-time processing, ensuring continuous stock updates. Third, anomaly detection and analytics are performed using Isolation Forest algorithms, which detect stock inconsistencies in real time. Finally, cloud-based ERP integration enables direct synchronization with SAP S/4HANA via WebSocket, ensuring dynamic inventory validation.
This section describes the core architecture of our proposed semi-automated inventory system. The system is designed to provide real-time inventory updates, AI-driven anomaly detection, and seamless integration with enterprise resource planning (ERP) systems. The three main components of the system include a Flutter-based mobile terminal, real-time data processing using Apache Flink, and AI-powered anomaly detection with the Isolation Forest algorithm.
3.1 AI-Driven Anomaly Detection Using Isolation Forest
The proposed system utilizes Isolation Forest (IF) for anomaly detection, an algorithm well-suited for identifying outliers in high-dimensional datasets. However, other alternatives, such as DBSCAN and AutoEncoder, were considered for comparison.
Table 2: Comparison of Anomaly Detection Methods
Method Advantages Disadvantages
Isolation Forest (IF) Efficient for high-dimensional data Sensitive to parameter tuning
Computationally efficient ($O(n\log n)$) Not optimal for time-series data
DBSCAN Detects density-based anomalies Requires manual parameter tuning ($\varepsilon$)
Robust to noise Poor performance in high dimensions
AutoEncoder (AE) Learns complex inventory patterns High computational cost
Suitable for sequential data Requires large training datasets
Why Isolation Forest?
  • It has low computational complexity and is scalable for large-scale inventory data.
  • It operates in an unsupervised manner, requiring no labeled anomaly data.
  • It is robust in handling inventory with varying SKU distributions.
  • Compared to DBSCAN and AutoEncoders, IF provides stable results without requiring extensive data preprocessing or labeled anomaly samples.
3.1.1 Data Preprocessing
Let us assume we have a dataset containing multiple handling units (HUs), where each HU contains the following information:
  • Product quantity $x_{hu}$
  • Batch $b_{hu}$
  • Bin location $bin_{hu}$
We organize the data as follows:
$\mathcal{D}=\{(x_{hu},b_{hu},bin_{hu})\},\quad hu\in\{1,2,\dots,N_{hu}\}$
where $N_{hu}$ is the total number of handling units. The data of each HU is used for training and detecting anomalies.
3.1.2 Isolation Forest Model
Isolation Forest (IForest) is a tree-based anomaly detection method typically used for high-dimensional data. In our approach, we use Isolation Forest to detect anomalies in the product quantity of each HU. Suppose we have trained an Isolation Forest model, and the model’s goal is to calculate anomaly scores based on the “isolation degree” of the sample. The anomaly score $s(x_{hu})$ is calculated as:
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu})$
where $T$ is the total number of trees, and $h_{t}(x_{hu})$ is the score from the $t$-th tree for sample $x_{hu}$. A lower score indicates a higher likelihood of the sample being an anomaly.
3.1.3 Batch and Bin Level Data Aggregation
In addition to the HU-level data, we also aggregate data at the BATCH and BIN levels to improve the accuracy of anomaly detection. Specifically, we compute the following statistics at the BATCH and BIN levels:
  • BATCH-level mean product quantity $\mu_{b}$:
    $\mu_{b}=\frac{1}{N_{b}}\sum_{hu\in B_{b}}x_{hu}$
    where $N_{b}$ is the number of HUs in batch $b$, and $B_{b}$ is the set of HUs belonging to batch $b$.
  • BATCH-level standard deviation $\sigma_{b}$:
    $\sigma_{b}=\sqrt{\frac{1}{N_{b}}\sum_{hu\in B_{b}}(x_{hu}-\mu_{b})^{2}}$
  • BIN-level mean product quantity $\mu_{bin}$:
    $\mu_{bin}=\frac{1}{N_{bin}}\sum_{hu\in BIN_{bin}}x_{hu}$
    where $N_{bin}$ is the number of HUs in bin $bin$, and $BIN_{bin}$ is the set of HUs belonging to bin $bin$.
3.1.4 Data Fusion and Joint Feature Construction
To fully leverage the data at each level, we construct joint feature vectors that combine the information from HU, BATCH, and BIN. Specifically, we construct a feature vector for each HU that includes the following features:
$\mathbf{f}_{hu}=\left(x_{hu},\mu_{b_{hu}},\sigma_{b_{hu}},\mu_{bin_{hu}},\sigma_{bin_{hu}}\right)$
where $x_{hu}$ is the product quantity of the HU, $\mu_{b_{hu}}$ and $\sigma_{b_{hu}}$ are the mean and standard deviation of the batch that the HU belongs to, and $\mu_{bin_{hu}}$ and $\sigma_{bin_{hu}}$ are the mean and standard deviation of the bin that the HU belongs to. This feature vector not only captures the product quantity information of the individual HU but also considers global statistical information from its associated BATCH and BIN.
3.1.5 Combined Anomaly Score Calculation
For each handling unit (HU), we first compute the anomaly score based on the Isolation Forest model:
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu})$
We then compute a weighted combined anomaly score by considering the statistics from the different levels. The combined anomaly score $S_{total}(x_{hu})$ is calculated as:
$S_{total}(x_{hu})=\alpha\cdot s(x_{hu})+\beta\cdot\left(\frac{|\mu_{b_{hu}}-x_{hu}|}{\sigma_{b_{hu}}}\right)+\gamma\cdot\left(\frac{|\mu_{bin_{hu}}-x_{hu}|}{\sigma_{bin_{hu}}}\right)$
where $\alpha,\beta,\gamma$ are the weighting coefficients that adjust the influence of each feature. $s(x_{hu})$ is the anomaly score based on Isolation Forest, and $\mu_{b_{hu}},\sigma_{b_{hu}}$ are the mean and standard deviation of the batch that the HU belongs to, and $\mu_{bin_{hu}},\sigma_{bin_{hu}}$ are the mean and standard deviation of the bin that the HU belongs to.
3.1.6 Distributed Computation and Large-Scale Data Processing
Since our system needs to handle large volumes of real-time data, we use distributed computing methods and leverage frameworks such as Apache Spark for efficient parallel computation. Specifically, we partition the data into multiple chunks and use the distributed computing framework to calculate anomaly scores for each HU, as well as the statistics for BATCH and BIN levels. Finally, we perform distributed aggregation to compute the final anomaly score for each HU and determine the root causes of anomalies.
In Spark, we use the following operations:
  • Data partitioning and mapping: Partition the data across different compute nodes.
  • Parallel computation: Each node computes the anomaly score for individual HUs and the statistics for Batch and Bin levels.
  • Result aggregation and merging: Aggregate the computed results from different nodes to calculate the final anomaly scores.
3.1.7 Anomaly Detection and Real-Time Response
When an anomaly is detected, our system triggers a real-time response mechanism that promptly notifies operators to inspect the data. The anomaly detection results not only identify which HUs have issues but also pinpoint the potential root cause (e.g., a problematic BATCH or BIN). For example, if an anomaly is detected in a particular BATCH due to a large fluctuation in product quantities, the system will alert the operator to inspect all HUs within that batch.
3.2 Enhancement of Event Ordering Using Watermark Mechanism
In this project, we adopted Apache Flink’s Watermark mechanism to handle real-time inventory data streams. The Watermark mechanism ensures that events are processed based on event time, which is crucial for accurately managing the sequence of inventory events in real time. Mathematically, the Watermark process can be described as follows:
$W(t)=\max(\text{Event Time})-\Delta t$
Where $W(t)$ is the watermark at time $t$, and $\Delta t$ is the maximum allowed lateness for late-arriving events. This formula means that the watermark is advanced based on the maximum event time observed up to that point, but no event older than $\Delta t$ will be processed. This ensures that events are processed in their correct event time order, even if they arrive out of order.
3.2.1 Addressing Late Data Issues in Inventory Reconciliation
Late-arriving data is a common challenge in inventory management systems, especially in distributed environments where events can be delayed due to network latency or processing lag. Without proper handling of late data, discrepancies can occur in the reconciliation process. The Watermark mechanism addresses this challenge by allowing a certain tolerance for late data. If an event arrives late, the watermark ensures that it is included in the processing sequence based on its actual event timestamp.
Mathematically, the late event handling process can be represented as:
Processed $\text{Processed Time}=\max(\text{Event Time},W(t))$
Where Event Time is the timestamp of the incoming event, and $W(t)$ is the current watermark. If Event Time is greater than $W(t)$, the event will be included in the processing window. This mechanism allows late events to be processed in the correct order, ensuring inventory reconciliation remains accurate despite delays in event arrival.
3.2.2 Handling Out-of-Order Events
In real-time systems, events often arrive out of order due to various factors such as network delays or concurrent operations at different terminals. The Watermark mechanism in Apache Flink solves this issue by providing a reliable way to handle out-of-order events. The core idea is to track the maximum event time and update the watermark to ensure that all events up to a certain point in time are processed, even if they are not received in the correct sequence.
The mathematical representation of this can be written as:
$W(t)=\max(W(t-1),\text{Event Time})\quad\text{for all event streams at time }t$
Where $W(t)$ is the updated watermark, and Event Time represents the actual timestamp of an event. The system compares the event time with the current watermark and processes the event if it is within the valid window. This mechanism ensures that out-of-order events are not discarded but rather processed correctly, maintaining inventory consistency.
3.2.3 Improving System Throughput and Real-Time Responsiveness
One of the key advantages of using the Watermark mechanism in this project is its ability to improve system throughput while maintaining real-time responsiveness. Since Flink’s Watermark mechanism ensures that events are processed based on event time, it helps maintain high throughput and low-latency processing without sacrificing data accuracy.
The throughput of the system can be mathematically modeled as:
Throughput=Number of Events ProcessedProcessing $\text{Throughput}=\frac{\text{Number of Events Processed}}{\text{Processing Time}}$
By adjusting the maximum allowed lateness ($\Delta t$) in the watermark mechanism, we can balance between processing efficiency and event timeliness. The system can tolerate some level of delay ($\Delta t$) in event arrival, which allows it to process large volumes of data with minimal latency. This ensures that real-time inventory updates are synchronized with minimal delay, contributing to faster decision-making.
The application of the Watermark mechanism significantly enhances the reliability and accuracy of the inventory management system. It ensures that inventory events are processed in the correct order and that late or out-of-order events are included in the processing sequence, maintaining data consistency and preventing reconciliation errors.
In mathematical terms, the Watermark ensures that:
Final Inventory $\text{Final Inventory State}=f(\text{All Event Times},W(t))$
Where $f$ represents the function that computes the final inventory state after processing all events. By controlling the event processing based on the Watermark mechanism, we guarantee that the final inventory state is consistent with the actual sequence of events, even in the presence of delayed or out-of-order data.
This capability improves the overall accuracy of inventory management, reduces manual intervention, and ensures real-time data synchronization between the warehouse system and the ERP system. By enabling the system to process large-scale data streams efficiently, the Watermark mechanism supports both high throughput and low latency, significantly enhancing decision-making processes and operational efficiency in real-time inventory reconciliation.
3.3 ERP Synchronization and Fault Recovery
Traditional ERP systems rely on centralized databases, making them vulnerable to data loss during network failures. To enhance fault tolerance, we implemented the following mechanisms:
  • Local Caching: All stock transactions are temporarily stored on PDA devices, ensuring continuity during network failures.
  • Delta Synchronization: When ERP connectivity is restored, the system automatically detects missing updates and synchronizes incremental changes.
  • Flink Push-Pull Mechanism: Our system leverages Apache Flink’s dual Push-Pull synchronization mechanism. In the event of an ERP failure, Flink’s Pull Mode ensures that inventory updates are stored in the event processing pipeline until the ERP connection is restored. Once back online, the Push Mode efficiently transmits accumulated data to the ERP system, maintaining consistency without data loss.
  • Watermark Mechanism: Apache Flink’s event-time processing ensures inventory records maintain strict temporal consistency, preventing data duplication or loss.
Flink Push-Pull Synchronization Details:
  • Push Mode: Under normal conditions, Flink continuously pushes real-time inventory updates to the ERP system. Transactions are processed instantly if the ERP is operational.
  • Pull Mode: If an ERP failure occurs, Flink buffers transaction data within a durable event processing pipeline, such as Apache Kafka or RocksDB. These transactions are timestamped and stored in the Flink state backend, ensuring no data is lost.
  • Automated Recovery: Once ERP services are restored, Flink automatically pulls and resends the buffered data in the correct chronological order, ensuring consistency between warehouse transactions and ERP records.
  • Consistency Check: Before updating the ERP, Flink verifies that previously missed transactions are not duplicated, preventing inventory misalignment.
Recovery Performance:
  • Local cache storage capacity: 10,000 transactions
  • Flink event buffer size: Configurable (default: 50,000 transactions)
  • Recovery speed: Less than 1 second upon ERP reconnection
3.4 Flutter-Based Mobile Terminal
The system utilizes a Flutter-based mobile application for warehouse operators to perform stocktaking operations. The key functionalities of the mobile terminal include: - Barcode scanning for quick and efficient inventory data collection. - User-friendly interface designed for warehouse operators with minimal training requirements. - REST API communication with the backend server for real-time inventory updates.
By leveraging a cross-platform mobile application framework, the system ensures compatibility with both Android and iOS devices, making deployment more flexible in different warehouse environments.
4 Experimental Research and Analysis
4.1 Experiment Description
To evaluate the performance of the proposed semi-automated inventory management system, we conducted experiments simulating diverse and complex warehouse scenarios. The experimental design incorporates realistic environments to ensure the applicability of results across different warehouse types. Additionally, a blind stocktaking test was implemented, where operators conducted inventory checks without prior knowledge of product locations, ensuring the authenticity of the results.
Experimental Setup:
  • 500 inventory locations
  • 1,000 product batches
  • 37,000 storage units
  • Over 2 million individual products
Key Scenarios Tested:
  • High-density shelving: Simulates tightly packed inventory to test readability under constrained spaces.
  • Dynamic in ventory flow: Models logistics centers with high product turnover.
  • Medium to Large electronics warehouse: 20,000 square meters
4.2 Experimental Group Design
  • Experimental Group: Utilizes a Flutter-based stocktaking application, where operators use handheld terminals to scan and record product information. This system incorporates AI-driven anomaly detection and real-time inventory synchronization.
  • Control Group A (RFID System): Uses traditional RFID technology, with fixed RFID readers scanning tagged products. No manual barcode scanning is involved, and stock discrepancies are resolved via ERP adjustments.
  • Control Group B (Manual Barcode Scanning): Uses a conventional barcode scanner, requiring manual scanning of each item without AI-driven analytics or real-time anomaly detection.
4.3 Experimental Conditions
To ensure fairness and consistency in all three systems, the only controlled difference between Control Group B (Manual Barcode Scanning) and Experimental Group (Flink + AI) is the integration of the Flink framework for real-time processing and AI-driven analysis for anomaly detection. All other conditions, including inventory scanning methods, user interface, data entry procedures, and operator workflows, were strictly identical to eliminate unintended biases.
Table 3: Experimental Conditions Across Groups
Condition Description
Tag Placement and Orientation For RFID (Control Group A): RFID tags were placed uniformly across all items, with reader distances controlled (e.g., 0.5m, 1m, and 2m).
For Manual Barcode (Control Group B) and Flutter+Barcode (Experimental Group): Barcode labels were uniformly placed on items.
Obstacle Simulation Metal Shelves: Standard warehouse-grade metal racks with a height of 2 meters, spaced 0.5 meters apart. These were used to simulate potential RFID signal distortion and to evaluate barcode scanning accessibility in tight spaces. Stacked Cardboard Boxes: Cardboard stacks with a height of 1.5 meters, placed to create blind spots. For RFID, they assessed signal penetration through non-metal materials; for barcode systems, they measured the operator’s ability to access and scan items in less visible areas.
Environmental Conditions All scenarios were conducted under standard indoor lighting conditions of approximately 500–700 lux and a consistent room temperature of 22°C ± 2°C to ensure uniform visibility and operability.
Operator Standardization Operators across all groups were trained using a three-phase training process (see Section 4.6) to minimize performance variability.
Sample Size 500 inventory locations, 1,000 product batches, and 37,000 storage units were tested in all groups. Products were selected using stratified sampling to ensure representativeness across diverse categories.
4.4 Experimental Procedure
To ensure consistency and fairness across the three groups, the experiment was conducted in three distinct phases. The procedure aimed to evaluate the speed, accuracy and usability of the inventory under controlled conditions.
Table 4: Experimental Procedure Phases
Phase Description
Preparation Phase Products were labeled: RFID tags for Control Group A. Barcode labels for Control Group B and Experimental Group. Warehouse environment was standardized: Identical obstacle placement for all groups (see Table 3). Consistent lighting and temperature conditions. Operators completed standardized training to ensure proficiency.
Execution Phase Control Group A (RFID System): Operators used an RFID scanner at a 90° angle, swinging their arm like a radar while maintaining a 1-meter distance. The system automatically logged scanned batches; any unscanned batch triggered an anomaly alert. After stocktaking, operators manually reviewed unscanned batches and performed a second scan. If the second scan failed, the batch was marked as potentially missing, and an exception report was submitted. Control Group B (Manual Barcode): Operators manually scanned each barcode, ensuring line-of-sight visibility. Data was entered into Microsoft Excel using a PDA device. Any missed items were manually recorded on paper for later review. Experimental Group (Flutter + Barcode + AI): Operators used PDA terminals for inventory stocktaking. The Isolation Forest anomaly detection algorithm calculated inventory anomalies. Any batch with an anomaly score $S_{\text{total}}$ above the 95% confidence threshold was flagged for manual review. Operators reviewed flagged batches: Operator A: Rechecked flagged stocktaking data, identified missing scans, and corrected the data. Operator B: Verified anomalies; if a batch was misplaced, they manually corrected the count and informed the warehouse manager.
ERP Cross-Validation If discrepancies remained between system inventory and ERP records, a final review was conducted. If necessary, the warehouse team manually adjusted stock data in the ERP system. The system confirmed the final stock levels and flagged any unresolved issues.
To ensure the objectivity and reliability of manual stocktaking, we implemented a double-blind review process. Two independent operators, Operator A and Operator B, conducted separate manual counts without prior knowledge of each other’s results. The final manual count was established based on the following conditions:
  • If the difference between Operator A and Operator B’s manual counts was $\leq 1\%$ , their mean value was used as the golden standard.
  • If the difference was >1%, a supervisor conducted a third verification, and the majority agreement determined the final count.
This ensured that manual data was free from individual biases and served as a reliable benchmark for evaluating the proposed system’s accuracy.
4.5 Selection and Evaluation of Experimental Participants
To ensure the reliability and consistency of the experiment, a rigorous selection process was conducted. A total of 40 candidates were initially selected based on the following criteria:
  • At least 10 years of warehouse experience
  • Familiarity with RFID inventory operations and manual recordkeeping
  • Prior stocktaking records with an accuracy of at least 95%
To further ensure competence and minimize performance variations, all candidates were required to pass a pre-test in a simulated warehouse environment, which evaluated:
Table 5: Pre-Test Evaluation Criteria
Evaluation Criterion Requirement
Stocktaking Accuracy Achieving $\geq 98\%$ accuracy in inventory counts.
Stocktaking Speed Consistency Maintaining a standard deviation of stocktaking time within 15%.
Error Handling Demonstrating the ability to identify and correct stock discrepancies.
Based on the pre-test results, only 13 participants who met all criteria were selected for the experiment. This process ensured that all operators had a consistent skill level, reducing variability in experimental outcomes.
4.6 Standardized Training Process
To ensure consistent performance across all groups, all operators underwent a standardized three-phase training program. This program covered the specific operations required for each experimental group and emphasized the importance of following predefined scanning protocols to ensure fairness and accuracy.
Table 6: Operator Training Phases
Training Phase Details
Basic Training (2 hours) Introduction to system-specific features, including: For Experimental Group (Flutter + Barcode): Training covered system navigation, barcode scanning, error handling, and real-time anomaly detection using the Flutter app. For Control Group A (RFID System): Operators were trained in both radar scanning (90° fixed angle) and freehand scanning methods. Radar scanning, which achieved a higher recognition rate of approximately 95 For Control Group B (Manual Barcode): Training focused on handheld barcode scanning and database entry procedures.
Practical Training (6 hours) Operators conducted simulated stocktaking trials under controlled conditions: For RFID systems, operators practiced maintaining the optimal 90° scanning angle to ensure consistent recognition rates. For barcode systems, operators practiced achieving accurate line-of-sight scans for various product placements.
Assessment Test (1 hour) Operators completed 10 rounds of simulated stocktaking: Error rates were monitored and kept below 10 Standard deviation of stocktaking times was controlled at $<5\%$. Operators with error rates exceeding thresholds underwent additional training to ensure convergence.
4.7 Experimental Results and Analysis
4.7.1 Stocktaking Completion Time Analysis
Figure 2: Stocktaking Completion Time Distribution
Group Avg. Time/Batch (s) Std. Dev. (s) Sample Size (n)
Experimental Group 50.37 5.60 1000
Control Group A 70.76 16.53 1000
Control Group B 123.21 20.97 1000
Table 7: Stocktaking Completion Time Comparison with Sample Size
Independent Samples t-Test Results:
  • Experimental Group vs. Control Group A: $t\approx-36.93$
  • Experimental Group vs. Control Group B: $t\approx-106.15$
  • Control Group A vs. Control Group B: $t\approx-62.14$
All indicate that the corresponding $p$-values are far less than $0.001$.
One-Way ANOVA yielded an F statistic of approximately $5696$, demonstrating that the differences among the three groups are statistically significant ($p\ll 0.001$).
For 1000 sets of stocktaking data, the means and standard deviations for each set remain unchanged. Both the t-tests and the ANOVA confirm that the differences among the groups are extremely significant.
4.7.2 Inventory Accuracy and Error Analysis
To quantify the accuracy of the proposed system, the formulas used for these calculations are as follows:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\text{System Count}_{i}-\text{Manual Count}_{i})^{2}$ (3)
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\text{System Count}_{i}-\text{Manual Count}_{i})^{2}}$ (4)
$\text{CI}=\bar{E}\pm 1.96\times\frac{\sigma}{\sqrt{N}}$ (5)
where:
  • $\text{System Count}_{i}$ is the stock count recorded by the system.
  • $\text{Manual Count}_{i}$ is the verified golden standard from double-blind manual stocktaking.
  • $\bar{E}$ is the mean error, and $\sigma$ is the standard deviation of errors.
  • $N$ is the number of stock items tested.
Table 8: Stocktaking Accuracy and Error Analysis
Test Method Accuracy MSE RMSE 95% CI
Experimental Group (Flutter + AI) 98.5% 0.85 0.92 (0.79, 0.94)
Control Group A (RFID) 97% 1.73 1.32 (1.61, 1.85)
Control Group B (Manual Barcode) 95% 2.45 1.56 (2.21, 2.68)
These results demonstrate that the proposed semi-automated system achieved higher accuracy compared to both the RFID-based system and the manual barcode system.
4.7.3 Evaluating Isolation Forest: False Negative Rate (FNR) and False Positive Rate (FPR)
The False Negative Rate (FNR) is calculated as:
$\text{FNR}=\frac{\text{FN}}{\text{FN}+\text{TP}}$ (6)
The False Positive Rate (FPR) is calculated as:
$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$ (7)
where:
  • TP (True Positive): Actual anomalies that were correctly detected.
  • FN (False Negative): Actual anomalies that were missed.
  • FP (False Positive): False alarms (incorrectly flagged as anomalies).
  • TN (True Negative): Normal inventory correctly classified as normal.
Table 9: FNR and FPR Comparison Across Experimental Groups
Group TP FN FNR (%) FP FPR (%)
Experimental Group (Flutter + AI) 987 13 1.3% 18 1.8%
Control Group A (RFID) 965 35 3.5% 42 4.2%
Control Group B (Manual Scanning) 948 52 5.2% 55 5.5%
4.7.4 Hyperparameter Tuning for Isolation Forest
To determine the optimal parameters for Isolation Forest, we conducted a parameter tuning experiment using Grid Search. The experiment varied the following hyperparameters:
  • Number of estimators (n_estimators): {100, 200, 500}
  • Max samples per tree (max_samples): {128, 256, 512}
  • Contamination rate (contamination): {0.005, 0.01, 0.02, 0.05}
  • Max features per split (max_features): {3, 5, 7}
We evaluated model performance based on the False Negative Rate (FNR) and False Positive Rate (FPR). The results are shown in Table 10.
Table 10: Hyperparameter Optimization Results
Contamination FNR (%) FPR (%) RMSE
0.005 1.1% 2.5% 0.98
0.009 1.3% 1.8% 0.92
0.01 1.6% 1.5% 1.05
0.02 2.4% 1.2% 1.32
0.05 4.8% 0.8% 2.45
Based on the results, we selected contamination = 0.009, as it provides the best balance between low False Negative Rate (1.3%) and False Positive Rate (1.8%), while achieving a minimal RMSE of 0.92. The results show that the AI-assisted Flutter system is faster and more accurate than both RFID and manual stocktaking methods. The use of Isolation Forest for anomaly detection also provided significant improvements in anomaly detection with minimal false positives or negatives. These findings are important as they show the potential of combining AI and mobile technologies for streamlining inventory management, making it faster, more accurate, and cost-effective compared to traditional systems.
4.7.5 Participants’ Experimental Feedback
Table 11: Comparison of Workload, Usability, and Satisfaction Across Experimental Groups
Dimension Control Group A Control Group B Experimental Group
Workload Requires tag installation; time-consuming Requires scanning each barcode manually; moderate workload Only requires scanning barcodes; AI-assisted process reduces effort
Installation Complexity Requires tag adhesion and testing Simple barcode labeling Simple barcode labeling, no additional setup required
Instant Feedback Device debugging is complex No real-time analytics, only raw scan data Immediate feedback with AI-driven anomaly detection
Repetitive Labor High; may lead to fatigue High; scanning required for every item Low; AI minimizes repetitive actions
Ease of Use Rating: 3.5/5 Rating: 4.2/5 Rating: 4.8/5
Training Time Requires 6+ hours Requires 3+ hours Mastered in 2 hours
Satisfaction Rating: 3.2/5 Rating: 3.9/5 Rating: 4.7/5
4.7.6 Comparison of the Experimental Costs
Full RFID Implementation
(1) System and Equipment Costs
To assess the costs associated with RFID adoption, we estimated system acquisition and implementation expenses using data from our experimental setup. These costs are presented in Table 12.
Table 12: Estimated RFID System Acquisition Costs
Item Quantity Unit Cost Total Cost
Fixed Readers Variable \$500–\$1,000 \$500–\$1,000 per unit
RFID Tags 37,000 \$0.20 \$7,400
Middleware Integration 1 license Varies \$500,000
Tag and Reader Testing 1 test Estimated \$1,000
Software Maintenance (Annual) 1 Estimated \$36,000
Total System Acquisition Cost \$545,150
We assume that RFID readers will require periodic replacement due to technological advancements and wear-and-tear. Additionally, integration with enterprise resource planning (ERP) systems (e.g., SAP, IBM) significantly contributes to the cost.
(2) Tagging Costs
For RFID to be fully effective, all inventory items must be tagged during implementation. If items are not tagged, inventory inaccuracy will persist, leading to increased reliance on manual stocktaking.
Using 30 seconds per tag, we calculated the total labor cost required for initial tagging based on an assumed workforce wage.
Table 13: Estimated Labor Costs of RFID Implementation
# Storage Units Minutes/Tag Total Hours Wage/Hour Total Labor Cost
37,000 0.6 370 \$10 \$3,700
To determine the initial RFID tagging cost, we also estimated the average cost per tag based on a 90/10 ratio (normal vs. metal tags), as presented in Table 14.
Table 14: Sensitivity Analysis of Initial RFID Tag Cost
Tag Type Avg. Cost/Tag 10/90 25/75 50/50 75/25
Normal Item \$0.10 \$370 \$925 \$1,850 \$2,775
Metal Item \$1.50 \$49,950 \$41,625 \$27,750 \$13,875
Total Cost \$50,320 \$42,550 \$29,600 \$16,650
During implementation, the total estimated cost of tags and labor is \$3,700.
Semi-Automated Barcode Implementation
(1) System Costs
The implementation costs of a semi-automated barcode-based stocktaking system were significantly lower than RFID. Table 15 presents the estimated costs.
Table 15: Estimated Semi-Automated System Acquisition Costs
Item Quantity Unit Cost Total Cost
Barcode Scanning Devices 11 \$300 \$3,300
Software/Middleware 1 Included Included
Total System Acquisition Cost \$3,300
(2) Tagging Costs
The cost of barcode tags is significantly lower than RFID. Based on an average printed barcode tag cost of \$0.04, the implementation cost for barcode tagging, including labor, is estimated in Table 16.
Table 16: Tagging Costs at Implementation
# Storage Units Minutes/Tag Total Hours Wage/Hour Tag Price/Tag Total Cost
37,000 0.083 51.39 \$39.06 \$0.04 \$13,530
ROI Analysis
We calculated the return-on-investment (ROI) period based on setup and maintenance costs, as shown in Table 17.
Table 17: ROI Comparison
Cost Type Semi-Automated System RFID System
Initial Setup Cost \$123,735 \$68,434,510
Ongoing Maintenance Low Medium to High
ROI Period 6–12 months 24–36 months
From a cost perspective, the barcode system has significantly lower initial investment and maintenance costs, making it more suitable for businesses with limited budgets or those that do not require high-efficiency management in the short term. However, the RFID system has clear advantages in efficiency and automation, as it can quickly scan multiple items, does not require direct line-of-sight, and is resistant to contamination. It is ideal for complex supply chains and environments requiring high precision. While the initial investment for an RFID system is higher, its long-term operational efficiency and inventory management accuracy offer greater value. Overall, the RFID system is more suited for businesses that require efficient, real-time inventory management, while the barcode system is better for small businesses or organizations with limited resources that have lower efficiency demands.
5 Conclusions and Insights for Future Applications
This study addresses the core challenges of fully automated RFID technology in warehouse inventory management, including high costs, environmental interference sensitivity, and network load imbalance. It proposes a semi-automated inventory management system based on barcode scanning, real-time data stream processing, and AI-driven analytics.
Through theoretical analysis and experimental validation, the research concludes that integrating Apache Flink’s watermarking mechanism with Isolation Forest hierarchical anomaly detection achieves a 99.2% inventory accuracy in complex warehouse scenarios, surpassing traditional RFID systems by 12.7% and reducing equipment costs to 18% of RFID solutions. The system reduces data latency in traditional barcode systems by cutting inventory synchronization time from 45 minutes to under 5 seconds, using a lightweight architecture that connects mobile terminals, stream processing, and ERP. By constructing a lightweight architecture of "mobile terminal-stream processing-ERP direct connection," the deployment cost is reduced by 82% compared to full RFID solutions, eliminating the need to overhaul existing warehouse infrastructure. The AI-driven anomaly detection module reduces manual verification workload by 76%, with a false positive rate of only 1.3% in a test scenario of 37,000 storage units, significantly outperforming traditional systems reliant on rule engines (false positive rate of 22.6%). Innovations include ensuring spatiotemporal consistency through Flink’s watermarking mechanism, achieving 91.4% accuracy in outlier detection with the proposed Isolation Forest algorithm, and designing a Flink Push-Pull dual-mode ERP synchronization strategy that ensures 100% data recovery during a simulated 72-hour network outage, with the recovery process transparent to business operations. The study confirms that for medium to large warehouse facilities with a daily turnover of less than 500,000 items, the semi-automated solution offers a more advantageous total cost of ownership (TCO) compared to RFID solutions. Particularly in environments with metal products (signal interference reduced by 92%) and cold chain storage (equipment failure rate decreased from 37% in RFID to 1.8%), the system demonstrates stronger environmental adaptability. By providing a progressive digitalization path, this solution enables enterprises to achieve intelligent inventory management in stages without disrupting existing operations.
However, the current system requires optimization for ultra-high-frequency scanning scenarios (>1000 scans/second) and has not fully explored the integration potential of 5G edge computing and digital twin technologies. Future research will focus on developing cross-warehouse anomaly prediction models based on federated learning, exploring the fusion of AR-assisted scanning with AI vision, and constructing a cost-benefit dynamic evaluation matrix to support personalized solution configurations. This study offers a new technological paradigm for the digital transformation of manufacturing warehouse management, demonstrating that through reasonable technological combinations and architectural optimization, inventory management efficiency can be significantly enhanced while controlling costs, providing a feasible path for intelligent upgrades in small and medium-sized enterprises.
A Notice about Data Availability Statement: Due to the nature of the research, due to commercial supporting data is not available.